You purchase 5 put option contracts with a premium of $3.86 and an exercise price of $80. If the stock price at expiration of the contracts is $86.27, what is your profit?
Answer #1

Exercise price = $80

Premium = $3.86

Stock price at the time of expiration = $ 86.27

No of put options = 5

Gross payoff = ($80 - $86.27) *5 = -6.27*5 = -$31.35

Premium = 5 * 3.86 = -$19.30

Net payoff = -$50.65

Therfore loss = -$50.65